OpenStudy (anonymous):

Please help me, I got B, am I right ? A system of equations is given below: y = 4x − 3 2x + 7y = 41 Which of the following steps could be used to solve by substitution? 2(y = 4x − 3) + 7y = 41 4x − 3 = 2x + 7y 2x + 7(4x − 3) = 41 2(y = 4x − 3)

OpenStudy (anonymous):

@Austin6i6

OpenStudy (anonymous):

which step do you think?

OpenStudy (anonymous):

it's not b

OpenStudy (anonymous):

so y=4x-3 in the first equation

OpenStudy (anonymous):

replace y in the second equation with (4x-3)

OpenStudy (anonymous):

So it would be c ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you. Can you help me with one more @10115658

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Solve the system of equations and choose the correct answer from the list of options. d + e = 15 −d + e = −5 Label the ordered pair as (d, e). (0, 0) (10, −5) (5, 10) (10, 5)

OpenStudy (anonymous):

first isolate e in the first equation

OpenStudy (anonymous):

can you do that?

OpenStudy (anonymous):

What do you mean by isolate

OpenStudy (anonymous):

get it by itself on one side of the equation

OpenStudy (anonymous):

e=...

OpenStudy (anonymous):

is it 5 ?

OpenStudy (anonymous):

so what you do is subtract d from both sides. e = 15 - d

OpenStudy (anonymous):

then you plug in (15-d) for e in the second equation

OpenStudy (anonymous):

is d 10 ?

OpenStudy (anonymous):

so now we have -2d + 15 = -5

OpenStudy (anonymous):

therefore d = 10

OpenStudy (anonymous):

now plug 10 in for d in the first equation 10 + e = 15

OpenStudy (anonymous):

therefore e=5

OpenStudy (anonymous):

So the answer would be (10,5 ) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

no problem